OpenStudy (anonymous):

In 1990, the life expectancy of females was 78.8 yr. In 2001, it was 79.8 yr. Let F(t) = life expectancy and t = the number of years since 1990. a) Find a linear function that fits the data

OpenStudy (anonymous):

t = 0, F = 78.8 t = 11, F = 79.8 slope = (79.8 - 78.8)/(11 - 0) = 1/11 so F = (1/11)t + 78.8

OpenStudy (anonymous):

The two data points that determine the line are (90, 78.8) and (101, 79.8), and we can use the two-point form F(t) - 78.8 = [ (79.8 - 78.8)/(101 - 90) ] (t - 90) F(t) = 78.8 + (1/11)(t - 90) F(t) = t/11 + 3884/55